OpenStudy (anonymous):

What is the solution of 3x + 5 = 2x – 7? the x-coordinate of the intersection point of the lines y = 3x + 5 and y = 2x – 7 the x-coordinates of the x-intercepts of the lines y = 3x + 5 and y = 2x – 7 the y-coordinate of the intersection point of the lines y = 3x + 5 and y = 2x – 7 the y-coordinates of the y-intercepts of the lines y = 3x + 5 and y = 2x – 7

OpenStudy (anonymous):

The solution of 3x+5=2x-7. x=-12 -31=-31